OpenStudy (anonymous):

find x- and y-intercepts 1. 5x+3=y

OpenStudy (anonymous):

to find the y intercept we would put 0 for the x, so then it would be 5(0)+3=y. so the y-intercept is y=3. to find the x intercept we would subtract the 3 to the other side , so it will be 5x=y-3. then we would use 0 as y. then the equations becomes 5x=-3, then we divide 5 then the x-intercept will be x=-3/5

OpenStudy (anonymous):

plzzz in a equation

OpenStudy (anonymous):

5x+3=y 5(0)+3=y 3=y for x intercept 5x+3=y -3 -3 5x=y-3 5x=0-3 5x=-3 divide 5 on both sides x=-3/5

OpenStudy (anonymous):

in y intercepts

OpenStudy (anonymous):

y=3